OpenStudy (anonymous):

The cost of producing x soccer balls in thousands of dollars is represented by h(x) = –5x + 6. The revenue is represented by k(x) = 9x – 2. Which expression represents the profit, (k – h)(x), of producing soccer balls?

OpenStudy (hybrik):

((9x-2)-(-5x+6))(x) = (9x-2+5x-6)(x) = (14x-8)(x)

OpenStudy (anonymous):

x= 4/7, 0?

OpenStudy (anonymous):

14x – 8 14x + 4 4x – 8 4x + 4

OpenStudy (anonymous):

those are my options @hybrik

OpenStudy (alekos):

my money is on the first one

OpenStudy (hybrik):

Like I said 14x-8, How did he get a medal?